OpenStudy (anonymous):

f(x) =\/x+9 ; g(x) = 8x - 13 Find f(g(x)).

OpenStudy (anonymous):

\/ 8x-13+9 ?

OpenStudy (anonymous):

so \/8x-4

OpenStudy (anonymous):

F(x) = x + 9 G(x) = 8x - 13 F(G(x)) = F(8x-13) = (8x-13) + 9 = 8x -13 +9 = 8x - 4

OpenStudy (anonymous):

so 2\/2x-1?

OpenStudy (anonymous):

Yeah, you're correct.

OpenStudy (anonymous):

4(2x-1)

OpenStudy (anonymous):

thanks. i have more..

OpenStudy (anonymous):

Send them.